Use the sum-to-product identities to rewrite the following expression as a product. cos(70°) + cos(100°)
Answer #1

using So 2 COS(A) + COS(B) = 2 2 cos(hte) cos(A-B Costo) + cos(100) = 2 cos (70+100 cos( 170-101 - 2 cos( 1970 -30) = 2008 (9